OpenStudy (anonymous):

*MEDAL* What is the integral of sin(x) from x=pi to 2pi?

OpenStudy (anonymous):

what is the anti derivative of sine?

OpenStudy (anonymous):

-cos(X)?

OpenStudy (mathmale):

Right. Know how to apply the limits of integration to this result?

OpenStudy (anonymous):

no, I have no idea

OpenStudy (anonymous):

plug in \(2\pi\) plug in \(\pi\) subtract the second value from the first

OpenStudy (anonymous):

whoa that was wrong!!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

make that \(-1-1=-2\)

OpenStudy (anonymous):

Thank you, I did not think it was that simple!

OpenStudy (anonymous):

that is why you love the fundamental theorem of calculus without it this would be almost impossible

OpenStudy (anonymous):

Okey-doke, thank you so much!

OpenStudy (anonymous):

yw